OpenStudy (anonymous):

75 degsin15deg+sin75degcos15deg=1 ?

OpenStudy (mathstudent55):

There's somethng missing in the beginning, right?

OpenStudy (anonymous):

cos

OpenStudy (mathstudent55):

sin(A + B) = sinAcosB + cosAsinB

OpenStudy (mathstudent55):

cos75sin15 + sin75cos15 = sin75cos15 + cos75sin15 = sin(75 + 15) = Can you continue now?

OpenStudy (anonymous):

yea i got it right its 1

OpenStudy (mathstudent55):

correct, since it sin(90) = 1

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm